How many grams of FeClwill be produced from 7.25 g of HCl in the following reaction?

FeS + 2HCl --> FeCl2 + H2S


1
Expert's answer
2022-02-15T12:51:02-0500

Moles of HCl = mass/MM

= 7.25g/(36.458g/mol)

= 0.199mol

Mole ratio of HCl:FeCl2 is 2:1

Moles of FeCl2 = (1/2)x 0.199mol = 0.099mol


Mass of FCl2 = Moles x MM

= 0.099mol x (126.751g/mol)

= 12.6g